# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v2

@@ Abstract
In this paper, we propose a new rotatable antenna (RA) model to improve the performance of wireless communication systems. Different from conventional fixed antennas, the proposed RA system can flexibly alter the three-dimensional (3D) boresight direction of each antenna independently by adjusting its deflection angles to achieve a desired array directional gain pattern. Specifically, we investigate an RA-enabled uplink communication system, where the receive beamforming and the deflection angles of all RAs at the base station (BS) are jointly optimized to maximize the minimum signal-to-interference-plus-noise ratio (SINR) among all the users. In the special single-user and free-space propagation setup, the optimal deflection angles of RAs are derived in closed form with the maximum-ratio combining (MRC) beamformer applied at the BS. Moreover, we analyze the asymptotic performance with an infinite number of antennas based on this solution, which theoretically proves that the RA system can achieve a higher array gain as compared to the fixed-antenna system. In the general multi-user and multi-path channel setup, we first propose an alternating optimization (AO) algorithm to alternately optimize the receive beamforming and the deflection angles of RAs in an iterative manner. Then, a two-stage algorithm that solves the formulated problem without the need for iteration is further proposed to reduce computational complexity. Simulation results are provided to validate our analytical results and demonstrate that the proposed RA system can significantly outperform other benchmark schemes.
@@ I Introduction
In the rapidly evolving landscape of global information and communications technology (ICT), the forthcoming sixth-generation (6G) wireless network is envisioned to support even more densely-connected users and devices across more diverse applications and services, thus demanding significantly higher performance targets compared to its preceding generations [1]. Undoubtedly, as one of the most critical technologies for the current fifth-generation (5G) mobile communication, multiple-input-multiple-output (MIMO) can dramatically enhance the transmission rate and reliability of wireless networks through beamforming and multiplexing via multiple antennas at the transceivers [2]. However, the channel capacity and spectrum efficiency achieved by conventional MIMO are insufficient to meet the stringent requirements of 6G in its new applications. To further improve the spatial resolution and degree of freedom (DoF), wireless networks tend to integrate drastically more antennas into arrays at the base station (BS), thereby advancing MIMO to massive MIMO, and even further to extremely large-scale MIMO [3, 4, 5, 6].
@@
Although larger-scale MIMO can offer substantial array and spatial multiplexing gains, it comes at the expense of much higher hardware costs and power consumption. Furthermore, increasing the number of antennas cannot fully exploit the spatial DoFs, since the traditional fixed antennas cannot change their positions or orientations flexibly once deployed. Recently, fluid antenna system (FAS) and movable antenna (MA) have been proposed as promising technologies to overcome this limitation and attracted growing attention in wireless communication [7, 8, 9, 10, 11]. Compared to fixed-antenna architecture, FAS/MA enables the local movement of antennas in a specified region through different antenna movement mechanisms, which can proactively reshape the wireless channels to more favorable conditions and thus achieve higher capacity without increasing the number of antennas. Furthermore, with such a new DoF offered at the physical layer, it has been validated that FAS/MA can achieve various significant performance advantages, including interference mitigation, flexible beamforming, and multiplexing enhancement [12, 13, 14]. By leveraging these capabilities of FAS/MA, substantial efforts have been devoted to integrating them with cutting-edge wireless technologies, such as integrated sensing and communications (ISAC) [15], unmanned aerial vehicle (UAV) communications [16], intelligent reflecting surface (IRS) [17, 18, 19], and over-the-air computation [20]. Nevertheless, while FAS/MA can bring numerous performance advantages, their practical implementation is highly constrained by the response time and/or movement speed of the antennas. Additionally, existing works on FAS/MA still face limitations in terms of spatial flexibility and performance enhancement since only the positions of antennas are adjusted while their orientations are fixed.
To fully exploit all six-dimensional (6D) spatial DoFs, 6D movable antenna (6DMA) has been recently proposed to flexibly adjust both the three-dimensional (3D) position and 3D rotation of distributed antennas/arrays [21, 22, 23]. Based on the long-term/statistical user channel distribution, the 6DMA-equipped transceiver can adaptively allocate its antenna resources to improve the array and spatial multiplexing gains. Additionally, 6DMA can achieve both array and geometric gains to enhance the performance of wireless sensing [24]. Although 6DMA provides a general model for position and rotation adjustable antennas, its implementation requires drastic changes of the current antenna architectures of existing BSs and thus may be practically cost-prohibitive. In addition, many new practical movement and rotation constraints should be considered in 6DMA systems [21, 22, 23], making the joint design of positions and rotations of all 6DMA arrays highly challenging and sophisticated.
@@
Motivated by the above, we propose in this paper a new antenna architecture, called rotatable antenna (RA), as a simplified implementation of 6DMA to improve the performance of wireless communications cost-effectively. In the RA system, the deflection angles of each directional antenna can be independently adjusted to change its 3D orientation/boresight, while its 3D position is kept constant to reduce the hardware cost and time/energy overhead associated with antenna position changes in 6DMA. Compared to the conventional fixed-antenna architecture, RA can enhance communication and sensing performance substantially by flexibly adjusting the antenna orientation/boresight within the full 3D space. In this way, RA provides a practical solution for enhancing array gains in desired directions to boost the transmit/received signal power, while reducing the radiation power in undesired directions to avoid information leakage and interference. Therefore, by strategically designing the beamforming and antenna orientation/boresight, the RA system can be deployed to further improve the array/multiplexing gain and enhance the sensing resolution/range in various applications, such as ISAC, massive machine-type communication (mMTC), simultaneous localization and mapping (SLAM), and so on.
@@
  • Since the adjustment of antenna orientation/boresight should consider both the propagation environment and the antenna directional gain pattern, we first introduce a pair of deflection angles to characterize the 3D boresight for each RA, and then present a new multi-path geometric near-field channel model related to the antenna deflection angles. Under this channel model, a minimum signal-to-interference-plus-noise ratio (SINR) maximization problem is formulated to jointly optimize the receive beamforming and the deflection angles of all RAs.
  • For the special single-user and free-space propagation setup with the maximum-ratio combining (MRC) beamformer applied at the BS, the optimal deflection angles of RAs are derived in closed form. Meanwhile, a closed-form expression and lower/upper bounds for the signal-to-noise ratio (SNR) under the uniform linear/planar array (ULA/UPA) setting are also derived, showing that the resultant SNR first increases linearly with the number of antennas and eventually converges to a certain limit. Additionally, an asymptotic analysis is provided for the case with an infinite number of antennas to theoretically demonstrate that with a larger range for antenna boresight adjustment, the proposed RA-enabled system can exploit more spatial DoFs to achieve a higher array gain.
  • For the general multi-user and multi-path channel setup, to tackle the formulated min-SINR maximization problem for balancing the array directional gains among different users over their multi-path channels, we first propose an alternating optimization (AO) algorithm that alternately optimizes the receive beamforming and the deflection angles of RAs in an iterative manner until convergence is attained. In particular, with the optimal minimum mean-square error (MMSE) beamformer applied at the BS, the subproblem that optimizes the deflection angles of RAs is first transformed into a pointing vector optimization problem and then solved by the successive convex approximation (SCA) technique. To reduce computational complexity, we further propose a two-stage algorithm that solves a weighted channel power gain maximization problem based on the zero-forcing (ZF) beamformer without the need for iteration.
  • Simulation results validate our theoretical analysis and demonstrate that the proposed RA-enabled system can significantly improve communication performance over various benchmark schemes. It is shown that with only a small range for antenna boresight adjustment, the RA-enabled system can reap considerable performance gains over the fixed-antenna system. Furthermore, the performance advantages of RAs in flexibly balancing the directional gain over multi-path channels become more pronounced with stronger antenna directivity.
@@
Notation: Upper-case and lower-case boldface letters denote matrices and column vectors, respectively. Superscripts $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ stand for the transpose, Hermitian transpose, and matrix inversion operations, respectively. The sets of $a\times b$ dimensional complex and real matrices are denoted by $\mathbb{C}^{a\times b}$ and $\mathbb{R}^{a\times b}$, respectively. $\mathcal{O}(\cdot)$ denotes the standard big-O notation. For a vector $\bm{x}$, $\|\bm{x}\|$ denotes its $\ell_{2}$-norm, $\mathrm{Re}\{\bm{x}\}$ denotes its real part, $\mathrm{diag}(\bm{x})$ returns a diagonal matrix with the elements in $\bm{x}$ on its main diagonal, and $[\bm{x}]_{a:b}$ denotes the subvector of $\bm{x}$ consisting of the elements from $a$ to $b$. For a matrix $\bm{X}$, $\mathrm{Tr}(\bm{X})$ and $\mathrm{rank}(\bm{X})$ denote its trace and rank, $[\bm{X}]_{a,b}$ denotes the $(a,b)$-th entry of matrix $\bm{X}$, $[\bm{X}]_{a:b,c:d}$ denotes the submatrix of $\bm{X}$ consisting of the elements located in rows $a$ to $b$ and columns $c$ to $d$, and $\bm{X}\succeq 0$ implies that $\bm{X}$ is positive semi-definite. $\bm{I}$ and $\bm{0}$ denote an identity matrix and an all-zero matrix, respectively, with appropriate dimensions. The distribution of a circularly symmetric complex Gaussian (CSCG) random vector with zero mean and covariance matrix $\bm{\Sigma}$ is denoted by $\mathcal{N}_{c}(\bm{0},\bm{\Sigma})$; and $\sim$ stands for “distributed as”.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider an RA-enabled uplink communication system, where $K$ users (each equipped with a single isotropic antenna) simultaneously transmit their signals in the same time-frequency resource block to a BS equipped with a UPA at a fixed position consisting of $N$ directional RAs. Without loss of generality, we assume that the UPA is placed on the $x$-$y$ plane of a 3D Cartesian coordinate system and centered at the origin with $N=N_{x}N_{y}$, where $N_{x}$ and $N_{y}$ denote the numbers of RAs along $x$- and $y$-axes, respectively. The separation between adjacent RAs is denoted by $d\leq\frac{\lambda}{2}$, where $\lambda$ denotes the signal wavelength, and the entire UPA size can be expressed as $N_{x}d\times N_{y}d$. The physical size of each RA element is denoted as $\sqrt{S}\times\sqrt{S}$ with $\sqrt{S}\leq d$, and we define $\xi\triangleq\frac{S}{d^{2}}\leq 1$ as the array occupation ratio of the effective antenna aperture to the overall UPA region. In general, for the hypothetical antenna element, we have $S=\frac{\lambda^{2}}{4\pi}$.
Assuming that both $N_{x}$ and $N_{y}$ are odd numbers for notational convenience, the reference position of RA $n$, which locates at the $n_{x}$-th column and $n_{y}$-th row on the UPA, can be expressed as
$\mathbf{w}_{n}\triangleq\mathbf{w}_{n_{x},n_{y}}\triangleq\mathbf{w}_{(n_{y}-1)N_{x}+n_{x}}=[n_{x}d,n_{y}d,0]^{T},$ (1)
where $n_{x}=0,\pm 1,\dots,\pm\frac{N_{x}-1}{2}$ and $n_{y}=0,\pm 1,\dots,\pm\frac{N_{y}-1}{2}$. Let $r_{k}$ denote the distance between the center of the UPA and user $k$, and the position of user $k$ is then denoted by $\mathbf{q}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Omega_{k}]^{T}$, with $\Phi_{k}\triangleq\operatorname{sin}\psi_{k}\operatorname{sin}\phi_{k}$, $\Psi_{k}\triangleq\operatorname{cos}\psi_{k}$, and $\Omega_{k}\triangleq\operatorname{sin}\psi_{k}\operatorname{cos}\phi_{k}$, where $\psi_{k}\in[0,\frac{\pi}{2}]$ and $\phi_{k}\in[0,2\pi]$ denote the zenith and azimuth angles of user $k$ with respect to the origin of the coordinate system, respectively. Accordingly, the distance between user $k$ and RA $n$ can be expressed as
$r_{k,n}$ $=\|\mathbf{q}_{k}-\mathbf{w}_{n}\|$
@@
where $\delta_{k}\triangleq\frac{d}{r_{k}}$. Note that $\delta_{k}\ll 1$ since the RA separation $d$ is typically on the order of wavelength in practice.
(a) Deflection angles
@@
Figure 2: Illustration of deflection angles and directional gain pattern of RA $n$.
II-A Antenna Boresight Adjustment
The original orientations/boresights of all RAs are assumed to be parallel to the $z$-axis, and the boresight of each RA can be independently adjusted in 3D direction mechanically and/or electrically through a common smart controller. Generally speaking, mechanical control allows for a larger adjustment range in terms of antenna orientation to equivalently and coarsely change boresight direction but its response latency cannot be ignored; while electrical control can quickly and finely adjust boresight direction (without changing antenna orientation) but has a relatively limited adjustment range.
As shown in Fig. 2(a), the 3D boresight adjustment of each RA can be described by a pair of deflection angles: the zenith and azimuth angles with respect to the $z$-axis. Specifically, for RA $n$, the zenith angle $\theta_{\mathrm{z},n}$ represents the angle between the boresight direction of RA $n$ and the $z$-axis, while the azimuth angle $\theta_{\mathrm{a},n}$ is the angle between the projection of the boresight direction of RA $n$ onto the $x$-$y$ plane and the $x$-axis. To characterize the 3D boresight direction of RA $n$, we define a pointing vector determined by the zenith angle $\theta_{\mathrm{z},n}$ and the azimuth angle $\theta_{\mathrm{a},n}$ of RA $n$ as
$\mathbf{f}(\bm{\theta}_{n})=\left[\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{cos}\theta_{\mathrm{a},n},\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{sin}\theta_{\mathrm{a},n},\operatorname{cos}\theta_{\mathrm{z},n}\right]^{T},$ (3)
where $\bm{\theta}_{n}\triangleq[\theta_{\mathrm{z},n},\theta_{\mathrm{a},n}]^{T}$ is defined as the deflection angle vector of RA $n$, and we have $\|\mathbf{f}(\bm{\theta}_{n})\|=1$ for normalization. To account for the antenna boresight adjustment range and to avoid antenna coupling between any two RAs, the zenith angle of each RA should be within a specific range:

### v4

@@ Abstract
In this paper, we propose a new rotatable antenna (RA) model to improve the performance of wireless communication systems. Different from conventional fixed antennas, the proposed RA system can flexibly and independently alter the three-dimensional (3D) boresight direction of each antenna to achieve a desired array directional gain pattern. Specifically, we investigate an RA-enabled uplink communication system, where the receive beamforming and the boresight directions of all RAs at the base station (BS) are jointly optimized to maximize the minimum signal-to-interference-plus-noise ratio (SINR) among all the users. In the special single-user and free-space propagation setup, the optimal boresight directions of RAs are derived in closed form with the maximum-ratio combining (MRC) beamformer applied at the BS. Moreover, we analyze the asymptotic performance with an infinite number of antennas based on this solution, which theoretically proves that the RA system can achieve a higher array gain than the fixed-antenna system. In the general multi-user and multipath channel setup, we first propose an alternating optimization (AO) algorithm to alternately optimize the receive beamforming and the boresight directions of RAs in an iterative manner. Then, a two-stage algorithm that solves the formulated problem without the need for iteration is proposed to further reduce computational complexity. Simulation results are provided to validate our analytical results and demonstrate that the proposed RA system can significantly improve the communication performance as compared to other benchmark schemes.
@@ I Introduction
In the rapidly evolving landscape of global information and communications technology (ICT), the forthcoming sixth-generation (6G) wireless network is envisioned to support even more densely connected users and devices. It will serve a wider range of applications and services, thus demanding significantly higher performance requirements compared to its preceding generations [1]. Undoubtedly, multiple-input multiple-output (MIMO) is one of the most critical technologies for the current fifth-generation (5G) mobile communication. MIMO can dramatically enhance the transmission rate and reliability of wireless networks. This is achieved through beamforming and multiplexing using multiple antennas at the transceivers [2]. However, the channel capacity and spectrum efficiency achieved by conventional MIMO are insufficient to meet the stringent requirements of 6G in its new applications. To further improve spatial resolution and degrees-of-freedom (DoFs), wireless networks tend to integrate drastically more antennas into arrays at the base stations (BSs), thereby evolving MIMO into massive MIMO, and ultimately into extremely large-scale MIMO [3, 4, 5, 6, 7].
@@
Although larger-scale MIMO can offer substantial array and spatial multiplexing gains, it comes at the expense of much higher hardware costs and power consumption. Furthermore, simply increasing the number of antennas cannot fully exploit the spatial DoFs, as traditional fixed antennas lack the flexibility to adjust their positions or orientations. Recently, fluid antenna system (FAS) and movable antenna (MA) have been proposed as promising technologies to overcome this limitation and have attracted growing attention in wireless communication [8, 9, 10]. Compared to the fixed-antenna architecture, FAS/MA enables the local movement of antennas in a specified region through different antenna movement mechanisms, which can proactively reshape the wireless channels to more favorable conditions and thus achieve higher capacity without increasing the number of antennas. Furthermore, with such a new DoF offered at the physical layer, it has been validated that FAS/MA can achieve various significant performance advantages, including interference mitigation, flexible beamforming, and multiplexing enhancement [11, 12, 13]. By leveraging these capabilities of FAS/MA, substantial efforts have been devoted to integrating them with cutting-edge wireless technologies, such as integrated sensing and communications (ISAC) [14], unmanned aerial vehicle (UAV) communications [15], intelligent reflecting surface (IRS) [16, 17, 18], and over-the-air computation [19]. Nevertheless, while FAS/MA can bring numerous performance advantages, their practical implementation is highly constrained by the response time and/or movement speed of the antennas. Additionally, existing works on FAS/MA still face limitations in terms of spatial flexibility and performance enhancement since only the positions of antennas are adjusted while their orientations are fixed. To fully exploit all six-dimensional (6D) spatial DoFs, 6D movable antenna (6DMA) has been recently proposed to flexibly adjust both the three-dimensional (3D) position and 3D rotation of distributed antennas/arrays [20, 21]. Based on the long-term/statistical user channel distribution, the 6DMA-equipped transceiver can adaptively allocate its antenna resources to improve the array and spatial multiplexing gains.
@@
Motivated by the above, we propose in this paper a new antenna architecture, called rotatable antenna (RA), as a simplified implementation of 6DMA to improve the performance of wireless communication cost-effectively. In the RA system, the 3D orientation/boresight of each directional antenna can be independently adjusted while its 3D position remains fixed. As such, unlike the translation movement in MA and 6DMA, which requires additional space such as sliding tracks and movable area, RA only requires local rotational adjustment, which can be more readily achieved through compact mechanical or integrated electronic design [22, 23]. This offers RA great scalability and compatibility with existing wireless systems. Compared to the conventional fixed-antenna architecture, RA can enhance communication and sensing performance substantially by flexibly adjusting the antenna orientation/boresight within the full 3D space. In this way, RA provides a practical solution for enhancing array gains in desired directions to boost the transmit/receive signal power, while reducing the radiation power in undesired directions to avoid information leakage and interference. Therefore, by strategically designing the beamforming and antenna orientation/boresight, the RA system can be deployed to further improve the array/multiplexing gain and enhance the sensing resolution/range in various applications, such as ISAC, massive machine-type communication (mMTC), simultaneous localization and mapping (SLAM), and others.
@@
  • Adjusting the antenna orientation/boresight requires judicious consideration of both the propagation environment and the antenna’s directional gain pattern. To facilitate this, we define a pointing vector to characterize the 3D orientation/boresight of each RA, and then construct a new multipath geometric near-field channel model based on the pointing vectors of all RAs. Under this channel model, we further formulate a minimum signal-to-interference-plus-noise ratio (SINR) maximization problem to jointly optimize the receive beamforming and the pointing vectors of all RAs.
  • For the special single-user and free-space propagation setup with the maximum-ratio combining (MRC) beamformer applied at the BS, we derive the optimal pointing vectors of RAs in closed form. Meanwhile, we also derive a closed-form expression and lower/upper bounds for the signal-to-noise ratio (SNR) under the uniform linear/planar array (ULA/UPA) setting, which show that the resultant SNR first increases linearly with the number of antennas and eventually converges to a certain limit. Additionally, we present an asymptotic analysis as the number of antennas goes to infinity, theoretically demonstrating that a wider rotational range for antenna boresight adjustment enables the proposed RA system to exploit more spatial DoFs, thereby achieving a higher array gain.
  • For the general multi-user and multipath channel setup, we formulate a minimum-SINR maximization problem to balance the array directional gains among users across their multipath channels. To address this problem, we first propose an alternating optimization (AO) algorithm that alternately optimizes the receive beamforming and the pointing vectors of RAs in an iterative manner until convergence is achieved. In particular, with the optimal minimum mean-square error (MMSE) beamformer applied at the BS, the subproblem that optimizes the pointing vectors of RAs is solved by the successive convex approximation (SCA) technique. To reduce computational complexity, we further propose a two-stage algorithm that solves a weighted channel power gain maximization problem based on the zero-forcing (ZF) beamformer without the need for iteration.
  • Simulation results validate our theoretical analysis and demonstrate that the proposed RA system can significantly improve communication performance over various benchmark schemes. It is shown that even with a small rotational range for antenna boresight adjustment, the RA system can reap considerable performance gains over the fixed-antenna system. Furthermore, the performance advantages of RA in flexibly balancing the directional gain over multipath channels become more pronounced with stronger antenna directivity.
@@
Notation: Upper-case and lower-case boldface letters denote matrices and column vectors, respectively. Superscripts $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ stand for the transpose, Hermitian transpose, and matrix inversion operations, respectively. The sets of $a\times b$ dimensional complex and real matrices are denoted by $\mathbb{C}^{a\times b}$ and $\mathbb{R}^{a\times b}$, respectively. $\mathcal{O}(\cdot)$ denotes the standard big-O notation. For a vector $\bm{x}$, $\|\bm{x}\|$ denotes its $\ell_{2}$-norm, $\mathrm{Re}\{\bm{x}\}$ denotes its real part, $\mathrm{diag}(\bm{x})$ returns a diagonal matrix with the elements in $\bm{x}$ on its main diagonal, and $[\bm{x}]_{a:b}$ denotes the subvector of $\bm{x}$ consisting of the elements from $a$ to $b$. For a matrix $\mathbf{X}$, $\mathrm{Tr}(\mathbf{X})$ and $\mathrm{rank}(\mathbf{X})$ denote its trace and rank, $[\mathbf{X}]_{a,b}$ denotes the $(a,b)$-th entry of matrix $\mathbf{X}$, $[\mathbf{X}]_{a:b,c:d}$ denotes the submatrix of $\mathbf{X}$ consisting of the elements located in rows $a$ to $b$ and columns $c$ to $d$, and $\mathbf{X}\succeq 0$ implies that $\mathbf{X}$ is positive semi-definite. $\mathbf{I}$ and $\mathbf{0}$ denote an identity matrix and an all-zero matrix, respectively, with appropriate dimensions. The distribution of a circularly symmetric complex Gaussian (CSCG) random vector with zero mean and covariance matrix $\bm{\Sigma}$ is denoted by $\mathcal{N}_{c}(\bm{0},\bm{\Sigma})$; and $\sim$ stands for “distributed as”.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider an RA-enabled uplink communication system, where $K$ users (each equipped with a single isotropic antenna) simultaneously transmit their signals in the same time-frequency resource block to a BS equipped with a UPA consisting of $N$ directional RAs. Without loss of generality, we assume that the UPA is placed on the $x$-$y$ plane of a 3D Cartesian coordinate system and centered at the origin with $N=N_{x}N_{y}$, where $N_{x}$ and $N_{y}$ denote the numbers of RAs along $x$- and $y$-axes, respectively. The separation between adjacent RAs is denoted by $\Delta$, and thus the entire UPA size can be expressed as $N_{x}\Delta\times N_{y}\Delta$. The physical size of each RA element is denoted by $\sqrt{A}\times\sqrt{A}$ with $\sqrt{A}\leq\Delta$, and we define $\xi\triangleq\frac{A}{\Delta^{2}}\leq 1$ as the array occupation ratio of the effective antenna aperture to the overall UPA region.
Assuming that both $N_{x}$ and $N_{y}$ are odd numbers for notational convenience, the reference position of RA $n$, which is located at the $n_{x}$-th column and $n_{y}$-th row on the UPA, can be expressed as
$\mathbf{w}_{n}\triangleq\mathbf{w}_{n_{x},n_{y}}\triangleq\mathbf{w}_{(n_{y}-1)N_{x}+n_{x}}=[n_{x}\Delta,n_{y}\Delta,0]^{T},$ (1)
where $n_{x}=0,\pm 1,\dots,\pm\frac{N_{x}-1}{2}$ and $n_{y}=0,\pm 1,\dots,\pm\frac{N_{y}-1}{2}$. Let $r_{k}$ denote the distance between the center of the UPA and user $k$ with $k=1,2,\dots,K$. Accordingly, the position of user $k$ is denoted by $\mathbf{u}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Omega_{k}]^{T}$, with $\Phi_{k}\triangleq\operatorname{sin}\psi_{k}\operatorname{sin}\phi_{k}$, $\Psi_{k}\triangleq\operatorname{cos}\psi_{k}$, and $\Omega_{k}\triangleq\operatorname{sin}\psi_{k}\operatorname{cos}\phi_{k}$, where $\psi_{k}\in[0,\pi]$ and $\phi_{k}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ denote the zenith and azimuth angles of user $k$ with respect to the origin of the coordinate system, respectively. Accordingly, the distance between user $k$ and RA $n$ can be expressed as
$r_{k,n}$ $=\|\mathbf{u}_{k}-\mathbf{w}_{n}\|$
@@
where $\delta_{k}\triangleq\frac{\Delta}{r_{k}}$ and $n=1,2,\dots,N$. Note that $\delta_{k}\ll 1$ since the RA separation $\Delta$ is typically on the order of wavelength in practice.
(a) Rotational angles
@@
Figure 2: Illustration of rotational angles and directional gain pattern of RA $n$.
II-A Antenna Boresight Rotation
The original orientations/boresights of all RAs are assumed to be parallel to the $z$-axis, and the boresight direction of each RA can be independently adjusted in 3D space by mechanical or electronic methods [22]. Specifically, mechanical control typically utilizes servo motors or micro-electromechanical systems (MEMS) to physically rotate the orientations of directional antennas, thereby adjusting their boresight directions [24]. In contrast, electronic control retains fixed antenna orientations while enabling boresight rotation through electronic techniques such as multi-feed switching or PIN diode biasing [25]. It is noted that mechanical control generally offers a wider range for boresight rotation, whereas electronic control provides much faster response and better compatibility with existing wireless systems.
As shown in Fig. 2(a), the 3D boresight direction of RA $n$ can be characterized by a pointing vector, defined as
$\vec{\mathbf{f}}_{n}=\left[f_{x,n},f_{y,n},f_{z,n}\right]^{T}\in\mathbb{R}^{3\times 1},$ (3)
where $f_{x,n}$, $f_{y,n}$, and $f_{z,n}$ are the projections of RA $n$’s pointing vector on the $x$-, $y$-, and $z$-axes, respectively. For the boresight rotation of RA $n$, we let $\theta_{\mathrm{z},n}$ represent its zenith angle (i.e., the angle between the boresight direction of RA $n$ and the $z$-axis) and $\theta_{\mathrm{a},n}$ represent its azimuth angle (i.e., the angle between the projection of the boresight direction of RA $n$ onto the $x$-$y$ plane and the $x$-axis). Accordingly, we have $f_{x,n}=\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{cos}\theta_{\mathrm{a},n}$, $f_{y,n}=\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{sin}\theta_{\mathrm{a},n}$ and $f_{z,n}=\operatorname{cos}\theta_{\mathrm{z},n}$. Furthermore, we have $\|\vec{\mathbf{f}}_{n}\|=1$ due to normalization. To account for practical rotational constraint and mitigate antenna coupling between any two RAs, the zenith angle of each RA should be confined to a specific range [26]: